# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ Appendix E $-\ell$ Converse Analysis for Near-Constant Weight Designs
$\frac{\partial g}{\partial d}=\log\Big{(}\frac{\sqrt{d^{2}-A^{2}}}{2\sqrt{\rho(1-\rho)}}\Big{)}-\log\Big{(}\frac{e^{\nu}-d}{e^{\nu}-1}\Big{)}.$ (222)
Note that $\frac{\partial g}{\partial d}$ has the same sign as $\frac{d^{2}-A^{2}}{4\rho(1-\rho)}-\frac{(e^{\nu}-d)^{2}}{(e^{\nu}-1)^{2}}$, which expands as
$g^{(1)}(d):=\Big{(}\frac{1}{4\rho(1-\rho)}-\frac{1}{(e^{\nu}-1)^{2}}\Big{)}d^{2}+\frac{2e^{\nu}}{(e^{\nu}-1)^{2}}d-\Big{(}\frac{A^{2}}{4\rho(1-\rho)}+\frac{e^{2\nu}}{(e^{\nu}-1)^{2}}\Big{)},$ (223)
where we only regard $d$ as the variable. We note $g^{(1)}(0)<0$ and perform some algebra to find that
$g^{(1)}(|A|)=-\Big{(}\frac{e^{\nu}-A}{e^{\nu}-1}\Big{)}^{2}\leq 0,\leavevmode\nobreak\ g^{(1)}(e^{\nu})=\frac{e^{2\nu}-A^{2}}{4\rho(1-\rho)}\geq 0.$ (224)
@@
  • When $4\rho(1-\rho)<(e^{\nu}-1)^{2}$, $g^{(1)}(d)$ has a unique zero in $d>0$. By (224) this zero falls in the range of feasible $d$ (i.e., $[|A|,e^{\nu}]$), so it is precisely the desired $d^{*}$, yielding
    $d^{*}$ $=\frac{-e^{\nu}+\sqrt{e^{2\nu}+(\frac{(e^{\nu}-1)^{2}}{4\rho(1-\rho)}-1)(\frac{A^{2}(e^{\nu}-1)^{2}}{4\rho(1-\rho)}+e^{2\nu})}}{\frac{(e^{\nu}-1)^{2}}{4\rho(1-\rho)}-1},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}4\rho(1-\rho)<(e^{\nu}-1)^{2}.$ (225)
  • When $4\rho(1-\rho)=(e^{\nu}-1)^{2}$, using (224) it is easy to see that $d^{*}$ is the unique zero of $g^{(1)}(d)$, and thus
    $d^{*}$ $=\frac{A^{2}+e^{2\nu}}{2e^{\nu}},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }4\rho(1-\rho)=(e^{\nu}-1)^{2}.$ (226)
  • When $4\rho(1-\rho)>(e^{\nu}-1)^{2}$, we can verify that the two zeros of $g^{(1)}(d)$ both fall in $(0,\infty)$. Combining with (224), we find that $d^{*}$ equals the smaller zero of $g^{(1)}(d)$:
    $d^{*}$ $=\frac{-e^{\nu}+\sqrt{e^{2\nu}+(\frac{(e^{\nu}-1)^{2}}{4\rho(1-\rho)}-1)(\frac{A^{2}(e^{\nu}-1)^{2}}{4\rho(1-\rho)}+e^{2\nu})}}{\frac{(e^{\nu}-1)^{2}}{4\rho(1-\rho)}-1},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}4\rho(1-\rho)>(e^{\nu}-1)^{2}.$ (227)
Therefore, $d^{*}$ defined in (221) is given by
@@
$\lim_{a\to 1}\frac{-e^{\nu}+\sqrt{e^{2\nu}+(a-1)(A^{2}a+e^{2\nu})}}{a-1}$ (229)
$=$ $\lim_{a\to 1}\frac{(a-1)(A^{2}a+e^{2\nu})}{(a-1)\big{(}\sqrt{e^{2\nu}+(a-1)(A^{2}a+e^{2\nu})}+e^{\nu}\big{)}}=\frac{A^{2}+e^{2\nu}}{2e^{\nu}}.$ (230)
@@
$\mathbb{P}\Big{(}G_{j,j^{\prime},1}>\big{(}\frac{1}{2}-\zeta\big{)}\frac{Cn\nu e^{-\nu}}{k}+\frac{1}{2}R_{j^{\prime}}\Big{)}\geq\exp\Big{(}-\frac{\nu ne^{-\nu}}{k}[f_{2}(C,\zeta,\rho,\nu)+o(1)]\Big{)}$ (232)
Recall that (C2) requires that (212) holds for some $j^{\prime}\in[p]\setminus s$, and the placements of the $p-k$ non-defective items are independent. Therefore, by arguments analogous to (75)–(81), we obtain the threshold
$n\leq\frac{(1-\eta_{2})\frac{1}{1-\theta}k\log\frac{p}{k}}{\nu e^{-\nu}\big{(}f_{2}(C,\zeta,\rho,\nu)+o(1)\big{)}}$ (233)
@@
To ensure the failure of MLE, we need (C1) and (C2) simultaneously hold for some feasible $(C,\zeta)$. By the above analyses, for given $(C,\zeta)$ we have (C1) and (C2) with $1-o(1)$ probability if both (202) and (233) hold, and by merging $\eta_{1},\eta_{2}$ into a single $\eta$, this can be written as
$n\leq\frac{(1-\eta)k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}}\frac{1}{\max\{\frac{1}{\theta}f_{1}(C,\zeta,\rho,\nu),f_{2}(C,\zeta,\rho,\nu)\}}$ (234)
for some $\eta>0$. Optimizing over $(C,\zeta)$ establishes the second term in (11).
@@ Appendix F $-\ell$ Achievability Analysis for Near-Constant Weight Designs
Recall that in this part of the analysis we only need to consider $\ell\in[1,\frac{k}{\log k}]$. We will establish the threshold for $n$ above which the restricted MLE decoder (29) has $o(1)$ probability of failing. As with the Bernoulli design, this is significantly more challenging than the converse. We suppose that the true defective set is $s=[k]$ without loss of generality.
@@
We first recap some notation that we also used when studying the Bernoulli design. For $\mathcal{J}\subset s$ with $|\mathcal{J}|=\ell$ we will use the notation $\mathcal{M}_{\mathcal{J}},\mathcal{M}_{\mathcal{J}0},\mathcal{M}_{\mathcal{J}1},\mathcal{N}_{0},\mathcal{N}_{00},\mathcal{N}_{01}$ and their corresponding cardinalities from Lemma 3. Note that conditioned on $\mathbf{X}_{s}$ and $\mathbf{Y}$, these index sets and their cardinalities are deterministic. For $\ell\in[1,\frac{k}{\log k}]$, we say that $(C,\zeta)\in[0,\infty)\times[0,1]$ is feasible if $\frac{Cn\nu e^{-\nu}\ell}{k}$ and $\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}$ are integers, subject to the restriction that $(C,\zeta)=(0,0)$ is the only feasible pair with $C=0$. For feasible $(C,\zeta)$, we define the set
with cardinality $k_{\ell,C,\zeta}=|\mathcal{K}_{\ell,C,\zeta}|$, as defined in (32). For given $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$, recall that $G_{\mathcal{J},\mathcal{J}^{\prime},1}$ denotes the number of tests in $\mathcal{N}_{01}\cup\mathcal{M}_{\mathcal{J}1}$ that contain some item from $\mathcal{J}^{\prime}$, and ${G}_{\mathcal{J},\mathcal{J}^{\prime},2}$ denotes the number of tests in $\mathcal{N}_{00}\cup\mathcal{M}_{\mathcal{J}0}$ that contain some item from $\mathcal{J}^{\prime}$.
@@
$G_{\mathcal{J},\mathcal{J}^{\prime},2}=M_{\mathcal{J},\mathcal{J^{\prime}}}+N_{0,\mathcal{J}^{\prime}}-G_{\mathcal{J},\mathcal{J}^{\prime},1}.$ (235)
@@
We seek to bound the probability that restricted MLE fails, considering each $\ell\in[1,\frac{k}{\log k}]$ separately. As before, we observe the following via Lemma 3(b): If restricted MLE fails and returns some $s^{\prime}$ with $|s\setminus s^{\prime}|\in[1,\frac{k}{\log k}]$, then for $\ell:=|s\setminus s^{\prime}|$, there exists some feasible $(C,\zeta)\in[0,\infty)\times[0,1]$ with respect to the specific $\ell$ such that (C1) and (C2) below simultaneously hold:
@@
    $G_{\mathcal{J},\mathcal{J}^{\prime},1}-G_{\mathcal{J},\mathcal{J}^{\prime},2}\geq(1-2\zeta)\frac{Cn\nu e^{-\nu}\ell}{k}.$ (236)
    By substituting (235), this can be equivalently formulated as
    $G_{\mathcal{J},\mathcal{J}^{\prime},1}\geq\Big{(}\frac{1}{2}-\zeta\Big{)}\frac{Cn\nu e^{-\nu}\ell}{k}+\frac{1}{2}\big{(}N_{0,\mathcal{J}^{\prime}}+M_{\mathcal{J},\mathcal{J}^{\prime}}\big{)}.$ (237)
In order to show that the overall failure probability is $o(1)$, we will first show that the probability of both (C1) and (C2) holding for some $(\ell,C,\zeta)$ is suitably small. Observe that for any $-\ell$ $\mathcal{J}\subset[p]$, we have $M_{\mathcal{J}}\leq\ell\Delta=\frac{\ell\nu n}{k}$, which implies that $k_{\ell,C,\zeta}=0$ holds for any $C>e^{\nu}$. Thus, (C1) does not hold when $C>e^{-\nu}$, so from now on we concentrate on a specific $\ell\in[1,\frac{k}{\log k}]$ and a given feasible $(C,\zeta)\in[0,e^{\nu}]\times[0,1]$.
@@
We first bound $\mathbbm{E}k_{\ell,C,\zeta}$ and then deduce the high-probability behaviour of $k_{\ell,C,\zeta}$ via Markov’s inequality. Since $k_{\ell,C,\zeta}=\sum_{\mathcal{J}\subset s,|\mathcal{J}|=\ell}\mathbbm{1}(M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k},\leavevmode\nobreak\ M_{\mathcal{J}0}=\zeta M_{\mathcal{J}})$, we have
@@
$\mathbbm{E}k_{\ell,C,\zeta}$ $=\binom{k}{\ell}\mathbb{P}\left(\text{for fixed }\mathcal{J}\subset s\text{ with }|\mathcal{J}|=\ell,\leavevmode\nobreak\ M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k},\leavevmode\nobreak\ M_{\mathcal{J}0}=\zeta M_{\mathcal{J}}\right)$ (238)
@@
$=\binom{k}{\ell}\mathbb{P}\left(M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k}\right)\mathbb{P}\left(\text{Bin}\Big{(}\frac{Cn\nu e^{-\nu}\ell}{k},\rho\Big{)}=\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}\right)$ (239)
since $\mathbf{X}_{s}$ (which determines $M_{\mathcal{J}}$) and $\mathbf{Z}$ (which determines $M_{\mathcal{J}0}$ under given $M_{\mathcal{J}}$) are independent.
@@
Bounding $\mathbb{P}(M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k})$: We make use of Lemma 7 in Appendix G, in which $\mathcal{W}^{(\mathcal{J})}$ is defined to index the tests that contain some item from $\mathcal{J}$, and its cardinality is defined as $W^{(\mathcal{J})}:=|\mathcal{W}^{(\mathcal{J})}|$. We thus have that $i\in\mathcal{M}_{\mathcal{J}}$ holds if and only if $i\in\mathcal{W}^{(\mathcal{J})}\setminus\mathcal{W}^{(s\setminus\mathcal{J})}$.
@@
Recall also the following notation introduced above (181): Given $s_{0}\subset s=[k]$, we let $\mathbf{T}_{s_{0}}$ be an unordered multi-set of length $|s_{0}|\Delta$ whose entries are in $[n]$, with the overall multi-set representing the $|s_{0}|\Delta$ placements from the items in $s_{0}$ in an unordered manner. It follows that $\mathcal{W}^{(\mathcal{J})}$ and $\mathcal{W}^{(s\setminus\mathcal{J})}$ are determined by the randomness of $\mathbf{T}_{\mathcal{J}}$ and $\mathbf{T}_{s\setminus\mathcal{J}}$, respectively. Given $\mathcal{J}$ satisfying $|\mathcal{J}|=\ell$ with $\frac{\ell}{k}=o(1)$, (319) and (321) in Lemma 7 give that for any $\delta>0$ we have $W^{(s\setminus\mathcal{J})}=\big{(}1-e^{-\nu}+\delta^{\prime}\sqrt{\frac{\ell}{k}}+o(1)\big{)}n$ for some $|\delta^{\prime}|<\delta$, with probability at least $1-2\exp(-2\nu^{-1}\delta^{2}n)$. Therefore, we set $\delta=\sqrt{C_{*}\frac{\nu\ell}{k}}$ (thus $\delta\sqrt{\frac{\ell}{k}}=o(1)$) for some sufficiently large absolute constant $C_{*}$ to obtain that the event $W^{(s\setminus\mathcal{J})}=(1-e^{-\nu}+o(1))n$, which implies $n-W^{(s\setminus\mathcal{J})}=ne^{-\nu}(1+o(1))$, holds with probability at least $1-2\exp(-\frac{C_{*}\ell n}{k})$. Thus, the event
$\mathscr{A}_{\mathcal{J}}=\Big{\{}n-W^{(s\setminus\mathcal{J})}=ne^{-\nu}(1+o(1))\Big{\}}$ (240)
@@
holds with probability at least $1-2\exp(-\frac{C_{*}\ell n}{k})$, where $C_{*}$ is a given constant that can be made arbitrarily large. Therefore, we can proceed as
$\mathbb{P}\Big{(}M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}$ $\leq\mathbb{P}\Big{(}M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k}\Big{|}\mathscr{A}_{\mathcal{J}}\Big{)}+\mathbb{P}\big{(}\mathscr{A}_{\mathcal{J}}^{c}\big{)}\leq\tilde{P}_{1}+2\exp\Big{(}-\frac{C_{*}\ell n}{k}\Big{)},$ (241)
@@
To bound $\tilde{P}_{1}$, we divide the $n$ tests into two parts $\mathcal{W}^{(s\setminus\mathcal{J})}$ and $[n]\setminus\mathcal{W}^{(s\setminus\mathcal{J})}$. Then, conditioning on $\mathbf{T}_{s\setminus\mathcal{J}}$ and using the randomness of $\mathbf{T}_{\mathcal{J}}$, we can identify $M_{\mathcal{J}}$ with the number of tests in $[n]\setminus\mathcal{W}^{(s\setminus\mathcal{J})}$ that contain item from $\mathcal{J}$. With this perspective, Lemma 8 in Appendix G states that
$\tilde{P}_{1}$ $\leq\exp\Big{(}-\frac{\ell\nu n}{k}\Big{[}D\Big{(}Ce^{-\nu}\big{\|}e^{-\nu}(1+o(1))\Big{)}+o(1)\Big{]}\Big{)}$ (242)
$=\exp\Big{(}-\frac{\ell\nu n}{k}\big{[}D(Ce^{-\nu}\|e^{-\nu})+o(1)\big{]}\Big{)}.$ (243)

### 23 Oct 2024 v2

@@ Appendix E $-\ell$ Converse Analysis for Near-Constant Weight Designs
$\frac{\partial g}{\partial d}=\log\Big{(}\frac{\sqrt{d^{2}-A^{2}}}{2\sqrt{\rho(1-\rho)}}\Big{)}-\log\Big{(}\frac{e^{\nu}-d}{e^{\nu}-1}\Big{)}.$ (E.47)
Note that $\frac{\partial g}{\partial d}$ has the same sign as $g^{(1)}(d):=\frac{d^{2}-A^{2}}{4\rho(1-\rho)}-\frac{(e^{\nu}-d)^{2}}{(e^{\nu}-1)^{2}}$, which can be expanded as
$g^{(1)}(d)=\Big{(}\frac{1}{4\rho(1-\rho)}-\frac{1}{(e^{\nu}-1)^{2}}\Big{)}d^{2}+\frac{2e^{\nu}}{(e^{\nu}-1)^{2}}d-\Big{(}\frac{A^{2}}{4\rho(1-\rho)}+\frac{e^{2\nu}}{(e^{\nu}-1)^{2}}\Big{)},$ (E.48)
where we only regard $d$ as the variable. We observe that $g^{(1)}(0)<0$, and perform some algebra to find that
$g^{(1)}(|A|)=-\Big{(}\frac{e^{\nu}-A}{e^{\nu}-1}\Big{)}^{2}\leq 0,\leavevmode\nobreak\ g^{(1)}(e^{\nu})=\frac{e^{2\nu}-A^{2}}{4\rho(1-\rho)}\geq 0.$ (E.49)
@@
  • When $4\rho(1-\rho)<(e^{\nu}-1)^{2}$, $g^{(1)}(d)$ has a unique zero in $d>0$. By (E.49), this zero falls in the range of feasible $d$ (i.e., $[|A|,e^{\nu}]$), so it is precisely the desired $d^{*}$, yielding
    $d^{*}$ $=\frac{-e^{\nu}+\sqrt{e^{2\nu}+(\frac{(e^{\nu}-1)^{2}}{4\rho(1-\rho)}-1)(\frac{A^{2}(e^{\nu}-1)^{2}}{4\rho(1-\rho)}+e^{2\nu})}}{\frac{(e^{\nu}-1)^{2}}{4\rho(1-\rho)}-1},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}4\rho(1-\rho)<(e^{\nu}-1)^{2}.$ (E.50)
  • When $4\rho(1-\rho)=(e^{\nu}-1)^{2}$, the right-hand side of (E.48) becomes linear in $d$, and solving gives
    $d^{*}$ $=\frac{A^{2}+e^{2\nu}}{2e^{\nu}},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }4\rho(1-\rho)=(e^{\nu}-1)^{2}.$ (E.51)
  • When $4\rho(1-\rho)>(e^{\nu}-1)^{2}$, we can verify that the two zeros of $g^{(1)}(d)$ both fall in $(0,\infty)$. Combining with (E.49), we find that $d^{*}$ equals the smaller zero of $g^{(1)}(d)$:
    $d^{*}$ $=\frac{-e^{\nu}+\sqrt{e^{2\nu}+(\frac{(e^{\nu}-1)^{2}}{4\rho(1-\rho)}-1)(\frac{A^{2}(e^{\nu}-1)^{2}}{4\rho(1-\rho)}+e^{2\nu})}}{\frac{(e^{\nu}-1)^{2}}{4\rho(1-\rho)}-1},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}4\rho(1-\rho)>(e^{\nu}-1)^{2}.$ (E.52)
Therefore, $d^{*}$ defined in (E.46) is given by
@@
$\lim_{a\to 1}\frac{-e^{\nu}+\sqrt{e^{2\nu}+(a-1)(A^{2}a+e^{2\nu})}}{a-1}$ (E.54)
$=$ $\lim_{a\to 1}\frac{(a-1)(A^{2}a+e^{2\nu})}{(a-1)\big{(}\sqrt{e^{2\nu}+(a-1)(A^{2}a+e^{2\nu})}+e^{\nu}\big{)}}=\frac{A^{2}+e^{2\nu}}{2e^{\nu}}.$ (E.55)
@@
$\mathbb{P}\Big{(}G_{j,j^{\prime},1}>\big{(}\frac{1}{2}-\zeta\big{)}\frac{Cn\nu e^{-\nu}}{k}+\frac{1}{2}R_{j^{\prime}}\Big{)}\geq\exp\Big{(}-\frac{\nu ne^{-\nu}}{k}[f_{2}(C,\zeta,\rho,\nu)+o(1)]\Big{)}.$ (E.57)
Recall that (C2) requires that (E.37) holds for some $j^{\prime}\in[p]\setminus s$, and the placements of the $p-k$ non-defective items are independent. Therefore, by arguments analogous to (C.39)–(C.45), we obtain the threshold
$n\leq\frac{(1-\eta_{2})\frac{1}{1-\theta}k\log\frac{p}{k}}{\nu e^{-\nu}\big{(}f_{2}(C,\zeta,\rho,\nu)+o(1)\big{)}}$ (E.58)
@@
Recall that for MLE to fail, it suffices that (C1) and (C2) simultaneously hold for some feasible pair $(C,\zeta)$. By the above analyses, for given $(C,\zeta)$ we have (C1) and (C2) with $1-o(1)$ probability if both (E.27) and (E.58) hold, and by merging $\eta_{1},\eta_{2}$ into a single parameter $\eta$, this can be written as
$n\leq\frac{(1-\eta)k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}}\frac{1}{\max\{\frac{1}{\theta}f_{1}(C,\zeta,\rho,\nu),f_{2}(C,\zeta,\rho,\nu)\}}$ (E.59)
for arbitrarily small $\eta>0$. Optimizing over $(C,\zeta)$ establishes the second term in (1.11).
@@ Appendix F $-\ell$ Achievability Analysis for Near-Constant Weight Designs
Recall that in this part of the analysis we only need to consider $\ell\in[1,\frac{k}{\log k}]$. We will establish the threshold for $n$ above which the restricted MLE decoder (B.1) has $o(1)$ probability of failing. As with the Bernoulli design, this is more challenging than the converse. We suppose that the true defective set is $s=[k]$ without loss of generality.
@@
We first recap some notation that we also used when studying the Bernoulli design. For $\mathcal{J}\subset s$ with $|\mathcal{J}|=\ell$ we will use the sets $\mathcal{M}_{\mathcal{J}},\mathcal{M}_{\mathcal{J}0},\mathcal{M}_{\mathcal{J}1},\mathcal{N}_{0},\mathcal{N}_{00},\mathcal{N}_{01}$ and their corresponding cardinalities from Lemma B.1. Conditioned on $\mathbf{X}_{s}$ and $\mathbf{Y}$, these sets and their cardinalities are deterministic. For $\ell\in[1,\frac{k}{\log k}]$, we say that $(C,\zeta)\in[0,\infty)\times[0,1]$ is feasible if $\frac{Cn\nu e^{-\nu}\ell}{k}$ and $\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}$ are integers, subject to the restriction that $(C,\zeta)=(0,0)$ is the only feasible pair with $C=0$. For a feasible pair $(C,\zeta)$, we define the set
$\mathcal{K}_{\ell,C,\zeta}=\Big{\{}\mathcal{J}\subset s,\,:\,|\mathcal{J}|=\ell,M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k},\leavevmode\nobreak\ M_{\mathcal{J}0}=\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}\Big{\}}$
with cardinality $k_{\ell,C,\zeta}=|\mathcal{K}_{\ell,C,\zeta}|$, as defined in (B.4). For given $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$, recall that $G_{\mathcal{J},\mathcal{J}^{\prime},1}$ denotes the number of tests in $\mathcal{N}_{01}\cup\mathcal{M}_{\mathcal{J}1}$ that contain at least one item from $\mathcal{J}^{\prime}$, and ${G}_{\mathcal{J},\mathcal{J}^{\prime},2}$ denotes the number of tests in $\mathcal{N}_{00}\cup\mathcal{M}_{\mathcal{J}0}$ that contain at least one item from $\mathcal{J}^{\prime}$.
@@
$G_{\mathcal{J},\mathcal{J}^{\prime},2}=M_{\mathcal{J},\mathcal{J^{\prime}}}+N_{0,\mathcal{J}^{\prime}}-G_{\mathcal{J},\mathcal{J}^{\prime},1}.$ (F.1)
@@
We seek to bound the probability that restricted MLE fails, first considering each $\ell\in[1,\frac{k}{\log k}]$ separately (before later applying a union bound). As before, we observe the following via Lemma B.1(b): If restricted MLE fails and returns some $s^{\prime}$ with $|s\setminus s^{\prime}|\in[1,\frac{k}{\log k}]$, then for $\ell:=|s\setminus s^{\prime}|$, there exists some feasible $(C,\zeta)\in[0,\infty)\times[0,1]$ with respect to the specific $\ell$ such that (C1) and (C2) below simultaneously hold:
@@
    $G_{\mathcal{J},\mathcal{J}^{\prime},1}-G_{\mathcal{J},\mathcal{J}^{\prime},2}\geq(1-2\zeta)\frac{Cn\nu e^{-\nu}\ell}{k}.$ (F.2)
    By substituting (F.1), this can be equivalently formulated as
    $G_{\mathcal{J},\mathcal{J}^{\prime},1}\geq\Big{(}\frac{1}{2}-\zeta\Big{)}\frac{Cn\nu e^{-\nu}\ell}{k}+\frac{1}{2}\big{(}N_{0,\mathcal{J}^{\prime}}+M_{\mathcal{J},\mathcal{J}^{\prime}}\big{)}.$ (F.3)
In order to show that the overall failure probability is $o(1)$, we will first show that the probability of both (C1) and (C2) holding for some $(\ell,C,\zeta)$ is suitably small. Observe that for any $-\ell$ $\mathcal{J}\subset[p]$, we have $M_{\mathcal{J}}\leq\ell\Delta=\frac{\ell\nu n}{k}$, which implies that $k_{\ell,C,\zeta}=0$ holds for any $C>e^{\nu}$. Thus, (C1) does not hold when $C>e^{-\nu}$, so from now on we concentrate on a specific $\ell\in[1,\frac{k}{\log k}]$ and a given feasible pair $(C,\zeta)\in[0,e^{\nu}]\times[0,1]$.
@@
We first bound $\mathbbm{E}k_{\ell,C,\zeta}$ and then deduce the high-probability behavior of $k_{\ell,C,\zeta}$ via Markov’s inequality. Since $k_{\ell,C,\zeta}=\sum_{\mathcal{J}\subset s,|\mathcal{J}|=\ell}\mathbbm{1}(M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k},\leavevmode\nobreak\ M_{\mathcal{J}0}=\zeta M_{\mathcal{J}})$, we have
@@
$\mathbbm{E}k_{\ell,C,\zeta}$ $=\binom{k}{\ell}\mathbb{P}\left(\text{for fixed }\mathcal{J}\subset s\text{ with }|\mathcal{J}|=\ell,\leavevmode\nobreak\ M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k},\leavevmode\nobreak\ M_{\mathcal{J}0}=\zeta M_{\mathcal{J}}\right)$ (F.4)
@@
$=\binom{k}{\ell}\mathbb{P}\left(M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k}\right)\mathbb{P}\left(\text{Bin}\Big{(}\frac{Cn\nu e^{-\nu}\ell}{k},\rho\Big{)}=\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}\right)$ (F.5)
since $\mathbf{X}_{s}$ (which determines $M_{\mathcal{J}}$) and $\mathbf{Z}$ (which determines $M_{\mathcal{J}0}$ given $M_{\mathcal{J}}$) are independent.
@@
Bounding $\mathbb{P}(M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k})$: We make use of Lemma G.3 in Appendix G, in which $\mathcal{W}^{(\mathcal{J})}$ is defined to index the tests that contain some item from $\mathcal{J}$, and its cardinality is defined as $W^{(\mathcal{J})}:=|\mathcal{W}^{(\mathcal{J})}|$. We thus have that $i\in\mathcal{M}_{\mathcal{J}}$ holds if and only if $i\in\mathcal{W}^{(\mathcal{J})}\setminus\mathcal{W}^{(s\setminus\mathcal{J})}$.
@@
Recall also the following notation introduced above (E.6): Given $s_{0}\subset s=[k]$, we let $\mathbf{T}_{s_{0}}$ be an unordered multi-set of length $|s_{0}|\Delta$ whose entries are in $[n]$, with the overall multi-set representing the $|s_{0}|\Delta$ placements from the items in $s_{0}$ in an unordered manner. It follows that $\mathcal{W}^{(\mathcal{J})}$ and $\mathcal{W}^{(s\setminus\mathcal{J})}$ are determined by the randomness of $\mathbf{T}_{\mathcal{J}}$ and $\mathbf{T}_{s\setminus\mathcal{J}}$, respectively. Given $\mathcal{J}$ satisfying $|\mathcal{J}|=\ell$ with $\frac{\ell}{k}=o(1)$, (G.12) and (G.14) in Lemma G.3 give that for any $\delta>0$ we have $W^{(s\setminus\mathcal{J})}=\big{(}1-e^{-\nu}+\delta^{\prime}\sqrt{\frac{\ell}{k}}+o(1)\big{)}n$ for some $|\delta^{\prime}|<\delta$, with probability at least $1-2\exp(-2\nu^{-1}\delta^{2}n)$. Therefore, we set $\delta=\sqrt{C_{*}\frac{\nu\ell}{k}}$ (thus $\delta\sqrt{\frac{\ell}{k}}=o(1)$) for some sufficiently large constant $C_{*}$ to obtain that the event $W^{(s\setminus\mathcal{J})}=(1-e^{-\nu}+o(1))n$, which implies $n-W^{(s\setminus\mathcal{J})}=ne^{-\nu}(1+o(1))$, holds with probability at least $1-2\exp(-\frac{C_{*}\ell n}{k})$. Thus, the event
$\mathscr{A}_{\mathcal{J}}=\Big{\{}n-W^{(s\setminus\mathcal{J})}=ne^{-\nu}(1+o(1))\Big{\}}$ (F.6)
@@
holds with probability at least $1-2\exp(-\frac{C_{*}\ell n}{k})$, where $C_{*}$ can be made arbitrarily large. Therefore, we can proceed as follows:
$\mathbb{P}\Big{(}M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}$ $\leq\mathbb{P}\Big{(}M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k}\Big{|}\mathscr{A}_{\mathcal{J}}\Big{)}+\mathbb{P}\big{(}\mathscr{A}_{\mathcal{J}}^{c}\big{)}\leq\tilde{P}_{1}+2\exp\Big{(}-\frac{C_{*}\ell n}{k}\Big{)},$ (F.7)
@@
To bound $\tilde{P}_{1}$, we divide the $n$ tests into two parts $\mathcal{W}^{(s\setminus\mathcal{J})}$ and $[n]\setminus\mathcal{W}^{(s\setminus\mathcal{J})}$. Then, conditioning on $\mathbf{T}_{s\setminus\mathcal{J}}$ and using the randomness of $\mathbf{T}_{\mathcal{J}}$, we can identify $M_{\mathcal{J}}$ with the number of tests in $[n]\setminus\mathcal{W}^{(s\setminus\mathcal{J})}$ that contain item from $\mathcal{J}$. With this perspective, Lemma G.4 in Appendix G states that
$\tilde{P}_{1}$ $\leq\exp\Big{(}-\frac{\ell\nu n}{k}\Big{[}D\Big{(}Ce^{-\nu}\big{\|}e^{-\nu}(1+o(1))\Big{)}+o(1)\Big{]}\Big{)}$ (F.8)
$=\exp\Big{(}-\frac{\ell\nu n}{k}\big{[}D(Ce^{-\nu}\|e^{-\nu})+o(1)\big{]}\Big{)}.$ (F.9)